# Tensor Product Attention Is All You Need

### v4

@@ D.2 Triton FlashTPA Decoding Kernel
7:$b\leftarrow p_{id_{B}}$; $h_{\text{start}}\leftarrow p_{id_{H}}\cdot B_{H}$
8:$m_{\text{block}\_\text{start}}\leftarrow p_{id_{M}}\cdot M_{\text{block}}$; $m_{\text{block}\_\text{end}}\leftarrow\min((p_{id_{M}}+1)\cdot M_{\text{block}},M)$
9$\triangleright$ $B_{H},B_{R},B_{D},B_{E}$ are tile sizes for dimensions H, R, D, E respectively.
@@
11$\triangleright$ Initialize accumulators for the head block
12:$\mathbf{o}_{\text{accum}}\leftarrow\mathbf{0}^{(E\times B_{H})}$; $\mathbf{m}_{\text{max}}\leftarrow-\infty^{(B_{H})}$; $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{0}^{(B_{H})}$; $c_{\text{scale}}\leftarrow s_{\text{total}}\cdot s_{Q}\cdot s_{K}$
@@
14$\triangleright$ Load query factors (fixed for this program as N=1)
@@
16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$triangleright Dimensions may be transposed after loading for matmul
@@
18$\triangleright$ Iterate over $M_{\text{chunk}}$-sized chunks within the K/V block
@@
23:    Load $\mathbf{a}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\mathbf{a}^{V}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\bm{b}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},D)$; $\bm{b}^{V}_{\text{chunk}}(E,M_{\text{curr}\_\text{chunk}})$triangleright Layouts optimized for memory access and matmuls
@@
26:    $S1_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{K}_{\text{chunk}},\mathbf{B}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},R_{Q})$
27:    $S2_{\text{chunk}}\leftarrow\text{MatMul}(S1_{\text{chunk}},\mathbf{A}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
28:    $S3_{\text{chunk}}\leftarrow S2_{\text{chunk}}\odot\mathbf{a}^{K}_{\text{chunk}}\cdot c_{\text{scale}}$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
@@
30:    $\mathbf{m}_{\text{max\_local}}\leftarrow\max_{axis=0}(S3_{\text{chunk}})$triangleright Shape: $(B_{H})$
@@
35:    $\mathbf{o}_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{V}_{\text{chunk}},\mathbf{p}_{\text{weighted\_av}})$triangleright Shape: $(E,B_{H})$
@@
39:    $\text{ratio}\leftarrow\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum}}$triangleright This is $\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum\_new}}$
@@
44$\triangleright$ Store partial results for this program’s (batch, head_block, M_block)
@@
46:$\mathbf{LSE}_{\text{val}}\leftarrow\log(\mathbf{s}_{\text{exp\_sum}})+\mathbf{m}_{\text{max}}$
@@
D.3 Additional Experimental Results
@@
Detailed Analysis of Figure 5 (Main Text; Embedding Dimension 2048): Figure 5 in the main paper depicts speed comparisons for an embedding dimension of 2048. The results indicate that FlashTPA (blue line) is highly competitive. Across all tested batch sizes (1 to 16) for $d_{\text{model}}=2048$:
@@
Appendix E Higher-Order Tensor Product Attention
@@
From a memory perspective during inference, higher-order TPA maintains the benefit of factorized KV caching. Only the constituent factors $\mathbf{a}_{K}(\mathbf{x}_{t}),\mathbf{b}_{K}(\mathbf{x}_{t}),\mathbf{c}_{K}(\mathbf{x}_{t})$ (and similarly for values) for each past token need to be stored. A trade-off arises between model capacity and the overhead of memory and computation. Higher-order tensor decompositions can provide additional flexibility and potentially increased capacity.
E.1 RoPE Compatibility in Higher-Order TPA
@@
Theorem E.1 (RoPE Compatibility in Higher-Order TPA).
@@
for $t\in\{1,\dots,T\}$ and $j\in\{1,\dots,d_{b}/2\}$. The transformation $\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$ operates on the $d_{h}$-dimensional vectorized features by post-multiplication. This structure of $\mathbf{T}_{t}$ ensures that the rotation effectively applied to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component (which is a column vector) corresponds to a pre-multiplication by $\mathbf{R}_{t}$, as detailed in the proof (Appendix F.2). This preserves the structure induced by $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ while rotating $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
@@
$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (E.1)
@@
Please see Appendix F.2 for the proof. For fourth-order or higher, this result still holds.
@@
0-shot 49.24 24.91 57.06 34.01 31.80 63.33 50.59 23.23 66.90 44.56
2-shot 53.37 25.34 48.78 34.00 29.20 62.79 52.33 26.41 75.30 45.28
Table 5: The evaluation results of medium models with third-order TPA pre-trained using FineWeb-Edu 100B dataset with lm-evaluation-harness. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
0-shot 56.19 29.44 58.36 43.74 33.80 67.95 52.80 24.07 77.20 49.28
2-shot 62.84 30.89 55.69 43.16 33.00 68.01 51.78 26.57 86.90 50.98
Appendix F Proofs of Theorems
F.1 Proof of Theorem 3.1
@@ Appendix F Proofs of Theorems
F.2 Proof of Theorem E.1
Theorem E.1 addresses the compatibility of RoPE with higher-order (specifically, 3rd-order) Tensor Product Attention. The theorem considers the query factorization:
@@
where $\mathbf{I}_{d_{c}}$ is the $d_{c}\times d_{c}$ identity matrix, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ is the standard RoPE block-diagonal matrix that pre-multiplies $d_{b}$-dimensional column vectors (as defined explicitly in the theorem statement in Section E).
@@
Appendix G More Related Work
@@
Appendix H More on Attention Mechanisms
H.1 Multi-Query Attention (MQA)
@@
H.2 Grouped Query Attention (GQA)
@@
H.3 Multi-head Latent Attention (MLA)
@@
$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (H.1)
$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (H.2)
where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix.
The term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding. However, as noted by [50], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations. RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section H.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation H.4). If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
[…]
Unlike Equation (H.2), acceleration by pre-computing the term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}\mathbf{T}_{t-s}(\bm{W}^{UK}_{i})^{\top}]$ is not possible because it depends on the relative position $(t-s)$ and thus varies for different $(t,s)$ pairs. To maintain RoPE compatibility while benefiting from compression, MLA introduces an additional, smaller key component $\mathbf{K}^{R}$ (and similarly $\mathbf{Q}^{R}$) to which RoPE is applied, while the main compressed components $\mathbf{K}^{C}$ and $\mathbf{V}^{C}$ (derived from $\mathbf{C}^{KV}$) remain RoPE-free. As we will demonstrate in Section 3.2 of the main paper, TPA offers a different approach to integrate RoPE efficiently with factorized attention through its tensor product formulation.
H.4 Multi-matrix Factorization Attention (MFA)
@@
H.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; 51) to encode positional information in the query/key vectors. Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head). $\mathbf{T}_{t}$ is a block-diagonal matrix composed of $d/2$ rotation blocks of the form $\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$ for $j\in\{1,\dots,d/2\}$. The frequencies $\{\theta_{j}\}$ are typically defined as $\theta_{j}=\text{base}^{-2j/d}$, with a common base like $10000$. If $\mathbf{q}_{t}\in\mathbb{R}^{d}$ is a query (or key) row vector for a specific head at position $t$, RoPE is applied as:
@@ H.5 Rotary Position Embedding (RoPE)
$\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (H.4)
which embeds relative positional information $(t-s)$ into the attention scores.
Appendix I More on TPA
@@
TPA with Non-contextual B. In Section A.1, we have introduced TPA with non-contextual A, where head-dimension factors $\mathbf{a}_{r}^{Q},\mathbf{a}_{r}^{K},\mathbf{a}_{r}^{V}\in\mathbb{R}^{h}$ are fixed. Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to the input token $\mathbf{x}_{t}$. The key tensor for token $t$, $\mathbf{K}_{t}\in\mathbb{R}^{h\times d_{h}}$, would then be constructed as:
@@
A similar formulation applies to values. This configuration might be effective if the fundamental token-level features (captured by $\mathbf{b}_{r}$) are relatively stable, while their combination across heads (captured by $\mathbf{a}_{r}(\mathbf{x}_{t})$) needs to adapt to the context. Performance comparisons for TPA with non-contextual A factors versus non-contextual B factors on small and medium-sized models are presented in Tables 6, 7, 8, and 9.
@@
[…]
@@
Appendix J More on Experiments
J.1 Experimental Settings

### v5

@@ B.1 Triton FlashTPA Decoding Kernel
[…]
@@
[…]
@@
[…]
@@
16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$ $\triangleright$ Dimensions may be transposed after loading for matmul
@@
[…]
@@
23:  Load $\mathbf{a}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\mathbf{a}^{V}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\bm{b}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},D)$; $\bm{b}^{V}_{\text{chunk}}(E,M_{\text{curr}\_\text{chunk}})$ $\triangleright$ Layouts optimized for memory access and matmuls
@@
26:  $S1_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{K}_{\text{chunk}},\mathbf{B}_{Q,\text{local}})$ $\triangleright$ Shape: $(M_{\text{curr}\_\text{chunk}},R_{Q})$
27:  $S2_{\text{chunk}}\leftarrow\text{MatMul}(S1_{\text{chunk}},\mathbf{A}_{Q,\text{local}})$ $\triangleright$ Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
28:  $S3_{\text{chunk}}\leftarrow S2_{\text{chunk}}\odot\mathbf{a}^{K}_{\text{chunk}}\cdot c_{\text{scale}}$ $\triangleright$ Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
@@
30:  $\mathbf{m}_{\text{max\_local}}\leftarrow\max_{axis=0}(S3_{\text{chunk}})$ $\triangleright$ Shape: $(B_{H})$
@@
35:  $\mathbf{o}_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{V}_{\text{chunk}},\mathbf{p}_{\text{weighted\_av}})$ $\triangleright$ Shape: $(E,B_{H})$
@@
39:  $\text{ratio}\leftarrow\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum}}$ $\triangleright$ This is $\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum\_new}}$
@@
[…]
@@
[…]
@@
B.2 Additional Experimental Results
@@
Detailed Analysis of Figure 5 (Embedding Dimension 2048): Figure 5 in the main paper depicts speed comparisons for an embedding dimension of 2048. The results indicate that FlashTPA (blue line) is highly competitive. Across all tested batch sizes (1 to 16) for $d_{\text{model}}=2048$:
@@
Appendix C Higher-Order Tensor Product Attention
@@
From a memory perspective, during inference, higher-order TPA maintains the benefit of factorized KV caching. Only the constituent factors $\mathbf{a}_{K}(\mathbf{x}_{t}),\mathbf{b}_{K}(\mathbf{x}_{t}),\mathbf{c}_{K}(\mathbf{x}_{t})$ (and similarly for values) for each past token need to be stored. A trade-off arises between model capacity and the overhead of memory and computation. Higher-order tensor decompositions can provide additional flexibility and potentially increased capacity.
C.1 RoPE Compatibility in Higher-Order TPA
@@
Theorem C.1 (RoPE Compatibility in Higher-Order TPA).
@@
for $t\in\{1,\dots,T\}$ and $j\in\{1,\dots,d_{b}/2\}$. The transformation $\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$ operates on the $d_{h}$-dimensional vectorized features by post-multiplication. This structure of $\mathbf{T}_{t}$ ensures that the rotation effectively applied to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component (which is a column vector) corresponds to a pre-multiplication by $\mathbf{R}_{t}$, as detailed in the proof (Appendix D.2). This preserves the structure induced by $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ while rotating $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
@@
$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr)=\operatorname{RoPE}_{t}\bigl(\mathbf{Q}_{t}\bigr),$ (C.1)
@@
Please see Appendix D.2 for the proof. For fourth-order or higher, this result still holds.
@@
0-shot 49.24 24.91 57.06 34.01 31.80 63.33 50.59 23.23 66.9 44.56
2-shot 53.37 25.34 48.78 34.00 29.20 62.79 52.33 26.41 75.3 45.28
Appendix D Proofs of Theorems
D.1 Proof of Theorem 3.1
@@ Appendix D Proofs of Theorems
D.2 Proof of Theorem C.1
Theorem C.1 addresses the compatibility of RoPE with higher-order (specifically, 3rd-order) Tensor Product Attention. The theorem considers the query factorization:
@@
where $\mathbf{I}_{d_{c}}$ is the $d_{c}\times d_{c}$ identity matrix, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ is the standard RoPE block-diagonal matrix that pre-multiplies $d_{b}$-dimensional column vectors (as defined explicitly in the theorem statement in Section C).
@@
Appendix E More Related Work
@@
Appendix F More on Attention Mechanisms
F.1 Multi-Query Attention (MQA)
@@
F.2 Grouped Query Attention (GQA)
@@
F.3 Multi-head Latent Attention (MLA)
@@
$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (F.1)
$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (F.2)
where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix. The term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding. However, as noted by [51], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations. RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section F.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation F.4). If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
$\begin{aligned} \mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}&=[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t}}^{\top}(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{s}}^{\top}(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]\\ &=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s}.\end{aligned}$ (F.3)
Unlike Equation (F.2), acceleration by pre-computing the term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}\mathbf{T}_{t-s}(\bm{W}^{UK}_{i})^{\top}]$ is not possible because it depends on the relative position (ts)(t-s) and thus varies for different $(t,s)$ pairs. To maintain RoPE compatibility while benefiting from compression, MLA introduces an additional, smaller key component $\mathbf{K}^{R}$ (and similarly $\mathbf{Q}^{R}$) to which RoPE is applied, while the main compressed components $\mathbf{K}^{C}$ and $\mathbf{V}^{C}$ (derived from $\mathbf{C}^{KV}$) remain RoPE-free. As we will demonstrate in Section 3.2 of the main paper, TPA offers a different approach to integrate RoPE efficiently with factorized attention through its tensor product formulation.
F.4 Multi-matrix Factorization Attention (MFA)
@@
F.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; 52) to encode positional information in the query/key vectors. Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head). $\mathbf{T}_{t}$ is a block-diagonal matrix composed of $d/2$ rotation blocks of the form $\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$ for $j\in\{1,\dots,d/2\}$. The frequencies $\{\theta_{j}\}$ are typically defined as $\theta_{j}=\text{base}^{-2j/d}$, with a common base like $10000$. If $\mathbf{q}_{t}\in\mathbb{R}^{d}$ is a query (or key) row vector for a specific head at position $t$, RoPE is applied as:
@@ F.5 Rotary Position Embedding (RoPE)
$\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (F.4)
which embeds relative positional information (ts)(t-s) into the attention scores.
Appendix G More on TPA
@@
TPA with Non-contextual B. In Section 4.1, we have introduced TPA with non-contextual A, where head-dimension factors $\mathbf{a}_{r}^{Q},\mathbf{a}_{r}^{K},\mathbf{a}_{r}^{V}\in\mathbb{R}^{h}$ are fixed. Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to the input token $\mathbf{x}_{t}$. The key tensor for token $t$, $\mathbf{K}_{t}\in\mathbb{R}^{h\times d_{h}}$, would then be constructed as:
@@
A similar formulation applies to values. This configuration might be effective if the fundamental token-level features (captured by $\mathbf{b}_{r}$) are relatively stable, while their combination across heads (captured by $\mathbf{a}_{r}(\mathbf{x}_{t})$) needs to adapt to the context. Performance comparisons for TPA with non-contextual A factors versus non-contextual B factors on small and medium-sized models are presented in Tables 5, 6, 7, and 8.
@@
$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\quad(\text{if contextual}),\quad\text{or}\quad\mathbf{b}^{K}_{r}=\mathbf{b}^{V}_{r}\quad(\text{if non-contextual}).$
@@
Appendix H More on Experiments
H.1 Experimental Settings